A local football stadium sells tickets to good seats for $5 each and poor seats for $2 each. On friday night there were 210 people in the stadium. The total ticket sale was $660. How many good and poor seats were sold?
Found 2 solutions by richwmiller, greenestamps:
Answer by richwmiller(17219)

total sold 210
total sales 660
product sold at 5 and 2
g +p =210,
5*g +2*p =660
g =210-p
We have two equations with two variables.
We will substitute to make one equation with one variable
5*(210-p )+2*p =660
1050-5p +2*p =660
-3*p =-390
p =130
g =210-p
g =80 at $5 and p =130 at $2
check
80+130=210
5*g +2*p =660
5*80+2*130=660
400+260=660
660=660
ok
codetickv

Answer by greenestamps(13200)



Here is a quick way to solve problems like this, if algebra is not required and an answer is to be obtained quickly -- as would be the case if a problem like this were on a competitive timed test.

All 210 tickets $2 each would bring in $420.
The actual total ticket sales was $660, which is $240 more than $420.
Each good ticket costs $3 more than each poor ticket.
The number of good tickets sold was $240/$3 = 80.